# Do Multimodal Large Language Models See Like Humans?

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have achieved impressive results on various vision tasks, leveraging recent advancements in large language models. However, a critical question remains unaddressed: do MLLMs perceive visual information similarly to humans? Current benchmarks lack the ability to evaluate MLLMs from this perspective. To address this challenge, we introduce HVSBench, a large-scale benchmark designed to assess the alignment between MLLMs and the human visual system (HVS) on fundamental vision tasks that mirror human vision. HVSBench curated over 85K multimodal samples, spanning 13 categories and 5 fields in HVS, including Prominence, Subitizing, Prioritizing, Free-Viewing, and Searching. Extensive experiments demonstrate the effectiveness of our benchmark in providing a comprehensive evaluation of MLLMs. Specifically, we evaluate 13 MLLMs, revealing that even the best models show significant room for improvement, with most achieving only moderate results. Our experiments reveal that HVSBench presents a new and significant challenge for cutting-edge MLLMs. We believe that HVSBench will facilitate research on human-aligned and explainable MLLMs, marking a key step in understanding how MLLMs perceive and process visual information.
@@
Figure 3: We are the first to systematically study and assess the alignment between MLLMs and HVS, by introducing diverse questions related to HVS. (a) We propose large-scale and comprehensive HVSBench, with a robust evaluation protocol. (b) We evaluate 13 leading MLLMs with HVSBench across 5 fields, revealing room for improvement and offering insights for developing HVS-aligned MLLMs. The shown predictions are from GPT-4o. For brevity, only 2 out of 10 human scanpaths are visualized. Questions and predictions are simplified.
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) have shown remarkable progress, achieving impressive performance across diverse vision-language tasks. These include applications such as image captioning [5], visual question answering [15, 30], document understanding [36], mathematic reasoning [29], and more. Such achievements highlight the capabilities of MLLMs in visual perception and vision-language interaction.  [^1] Equal contributions. Project page: https://jiaying.link/HVSBench/
@@
Despite the impressive performance of MLLMs on vision tasks, we have limited understanding of why they perform well. It has been demonstrated in previous studies that principles inspired by the HVS and grounded in cognitive science [11], play a vital role in enhancing the performance of backbone models, such as attention-based architectures [41]. However, discrepancies remain in how MLLMs and humans perceive visual information. Human attention is based on innate and learned saliency, while MLLMs often perceive images as arrays of pixel values or feature embeddings. Human visual attention is sequential, adjusting based on context and prior knowledge, while MLLMs process input statically or through fixed-length attention. Human attention can be dynamically influenced and guided by goals, while MLLMs lack the cognition and the ability to “refocus” dynamically, relying purely on trained associations. Given the importance of HVS-aligned designs in enhancing performance and enabling applications, like content generation, assistive tools, etc., the question of whether and to what extent existing MLLMs align with HVS remains a critical area for research, especially considering the current disparities.
@@
This brings us to a fundamental question: Do MLLMs perceive the world in the same way as humans do? More specifically, do MLLMs fixate on similar regions of interest within an image or follow a similar temporal order as the HVS when perceiving an image? Humans can effortlessly identify objects that capture their attention and perform visual searches based on context. However, in real-world scenarios, MLLMs often perceive visual information differently. For instance, in the first image of Fig. 3, if we ask, “Which object is the most visually prominent?”, most people would choose the left person, while MLLMs identify the van. Similarly, when searching for a chair in the second image, humans tend to first identify related objects (e.g., a table) and use contextual cues to help locate the chair (e.g., chairs are often near tables), while MLLMs may point to irrelevant areas. Therefore, we propose using various questions to assess MLLMs’ alignment across different HVS fields.
However, there has been limited research on how current MLLMs align with the HVS. Evaluating this alignment presents significant challenges. On the one hand, existing public vision-language datasets are primarily designed to assess model’s performance on specific tasks, offering little insight into its alignment with HVS. On the other hand, traditional HVS research has largely focused on low-level, vision-only domains [32], often evaluating models based on masks or heatmaps, making it difficult to assess MLLMs where the primary output is text rather than visual data.
@@
We thoroughly evaluate 13 well-known SOTA MLLMs on HVSBench, spanning diverse architectures and model scales. This not only provides a direct comparison among these models across multiple aspects of the HVS, but also highlights the significant gap between current MLLMs and humans. Furthermore, our findings reveal critical insights for future improvements: aligning MLLMs with the HVS cannot be achieved merely by incorporating external knowledge of related cues and priors, or by integrating human-generated captions and summaries. Additionally, we design a prominence enhancement to demonstrate how human-aligned MLLMs can improve downstream applications. Our contributions are summarized as follows:
@@
    We conduct a comprehensive evaluation of 13 popular MLLMs using HVSBench. Additionally, we provide new insights and techniques for developing more HVS-aligned and explainable MLLMs, including a prominence enhancement method that shows how HVS-aligned MLLMs improve downstream applications.
@@ 2 Related Work
Benchmarks for MLLMs. Existing benchmarks focus on assessing MLLMs’ capabilities in understanding and reasoning across modalities. A wide range of benchmarks, such as MMBench [27] and SEEDBench [22], have been proposed to evaluate general multimodal capabilities, including visual perception, reasoning, and comprehension. These benchmarks span diverse tasks, including document understanding [33], visual question answering [31], hallucination detection [25] and mathematical reasoning [29].
The introduction of HVSBench addresses this gap, focusing on human-centric evaluation to reveal the secret of how MLLMs perceive and process visual information.
@@ 3 HVSBench
In this section, we present our HVSBench in detail. We begin by defining the five distinct fields of HVS, followed by a description of multimodal QA generation, and automatic standardization and evaluation protocol. Fig. 4 shows some examples in our HVSBench. The Sec. A and Sec. D in supplementary materials provide the rationale behind selecting fields, their corresponding source datasets, additional sample questions and the results for each question type.
@@ 3.1 Preliminaries
Subitizing. Subitizing [48] is to quickly and accurately perceive the number of visually prominent objects in a scene. Compared to Prominence, it requires simultaneous attention to multiple elements. It is crucial in real-world scenarios where humans need to quickly estimate the number of prominent items, such as in navigation or crowd analysis.
Prioritizing. Prioritizing in HVS enables humans to rank objects within a scene based on their perceptual saliency [10]. It better captures the dynamic nature of HVS, i.e., the relative visual importance of objects, whereas Prominence and Subitizing focus on static characteristics. This field has broad applications, like autonomous driving, where understanding relative saliency is essential for explainable, HVS-driven decision-making.
Free-Viewing. Free-viewing is crucial for both behavioral vision in the HVS and computer vision applications. It focuses on the sequence of human gaze behavior in a task-free context [44]. This behavior is bottom-up, driven solely by the intrinsic processing of visual stimuli. Since free-viewing closely mirrors human gaze behavior, evaluating a model’s ability to predict both the order and duration of fixations, as well as the prediction of a full scanpath, provides insight into its alignment with the HVS.
@@
For each field, we select a curated collection of diverse, high-quality source datasets with human annotations, to create large-scale and high-quality multimodal QAs.
@@ 3.2 Benchmark Curation
(1) Template-Based Construction: For most questions, we generate answer options directly from ground-truth annotations. For example, in fields like Prominence, the options are derived from bounding boxes or other directly annotated features. To ensure diversity in the questions, we design multiple templates for each question type, following the methodology outlined in [26]. These templates provide variation while maintaining the focus of the field. For instance, in Prominence, we use templates such as:
@@ 3.3 Automatic Standardization and Evaluation
For the Prominence field, the automatic standardization cleans and normalizes predictions into consistent main choice labels (e.g., ‘A’, ‘C’, etc.) from various input formats, using regex to detect explicit or implied labels in diverse phrasing. For Subitizing, the automatic standardization processes predictions to extract integer values, including direct numbers, text-based numbers (e.g., “three”), or implied quantities. When no valid number is found, it defaults to the average ground truth value in 3.2116. For Prioritizing, except for choice answers that are processed the same as in the Prominence field, our automatic standardization detects and normalizes answers of sequence (sorting) via order-related phrasing through regex patterns. In Free-viewing and Searching fields, The automatic standardization involves extracting and normalizing scanpath data (X, Y for coordinates and T for durations) from diverse input formats, including text, JSON, and various phrasing styles. It also addresses irregularities like mismatched list lengths and incomplete scanpaths. Refer to Sec. B in supplementary for predictions and their corresponding standardized outputs in each field.
We illustrate the superiority of our automatic standardization over LLM-based matching in Fig. 5. In our experiments, GPT-4 was prompted with the full question, the full prediction, and the text prompt: “Please extract the prediction with the correct format.” Refer to Sec. C in supplementary for detailed inputs, experimental settings, and pseudo-code for our automatic standardization.
@@
Evaluation Metrics. In the Prominence field in which are choice questions, the accuracy can be calculated in $\text{Acc}=\frac{\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i})}{N}$, where $y_{i}$ is the correct answer for the $i$-th question, $\hat{y}_{i}$ is the prediction for the $i$-th question., $\mathbb{I}(\hat{y}_{i}=y_{i})$ is the indicator function that equals 1 if $\hat{y}_{i}=y_{i}$ and 0 otherwise, and $N$ is total number of predictions. For the Subitizing field, we adopt widely used metrics, including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), as well as Exact Match Accuracy (Acc). The MAE is calculated by $\text{MAE}(GT,pred)=\frac{1}{N}\sum_{i=1}^{N}\lvert\text{GT}_{i}-\text{pred}_{i}\rvert$ and RMSE is $\text{RMSE}(GT,pred)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\text{GT}_{i}-\text{pred}_{i})^{2}}$. The Acc is expressed in the same way as defined before. For Prioritizing field, the answer types include single-choice and sequence (sorting). The single choices are evaluated via accuracy. Then, to evaluate whether the predicted order of prioritized items matches the ground truth exactly, Exact Match Accuracy (Acc) is also used. In Free-viewing and Searching fields, we employ the widely adaptive MultiMatch [17] similarity to evaluate generated scanpaths. It measures scanpath similarity across multiple dimensions: shape, direction, length, and position. In our main results, we focus on the direction (”M-Dir”) and position (”M-Pos”) dimensions for simplicity. Following standard procedures [43], predictions are cropped to a maximum length of 6, and the final scores with each prediction being the mean of the MultiMatch calculated across all 10 GT scanpaths in the image.
@@ 4.1 Evaluation Details
We use different approaches for each field and specialized metrics to evaluate the performance of random guesses. In the Subitizing section, the accuracy metric is computed using random sampling based on the frequency distribution of subitized object counts. For MAE and RMSE, the average number of salient objects across all samples is used as the predicted value. For metrics related to MultiMatch (i.e., M-Dir and M-Pos), following [43], we randomly select a human scanpath in a different input image as its output and calculate the average performance based on it accordingly.
We select recent state-of-the-art open-source MLLMs for validation, including MiniGPT4-v2 [50], DeepSeek-VL [28], Idefics-series [21, 20] LLaVA-Next [26], LLaVA-OneVision (LLaVA-OV) [23], mPLUG-Owl3 [46], Qwen2-VL (Qwen2) [42], InternVL2.0 [7] and MiniCPM-V 2.6 [45]. We utilize the $\sim$7B-scale versions of these models, a widely adopted choice for balancing efficiency and effectiveness. We also include GPT4-o [1] and Gemini-1.5 Flash [37], two representative proprietary MLLMs, as key benchmarks in our evaluation. All experiments are conducted using VLMEvalKit [12] on the same platform to ensure consistency and fairness.
@@ 4.2 Main Results on HVSBench
Quantitative Evaluation. Table 3 presents the evaluation results on HVSBench, highlighting that current open-source and proprietary MLLMs still underperform in aligning with HVS. For example, GPT4-o falls short in several metrics, such as Prominence, Prioritizing, and Searching accuracy, where it performs worse than several open-source models such as Qwen2 [42]. Notably, Qwen2 [42] achieves the best overall performance, with LLaVA-OV [23] following closely behind. Both models even surpass GPT4-o overall, demonstrating the potential of recent open-source MLLMs.
@@
However, almost all methods perform poorly on scanpath similarity metrics (“M-Dir” and “M-Pos” in Free-viewing (Q9) and Searching (Q12)), underperforming random guessing, except for Qwen2 [42] in searching. These results suggest significant room for scanpath prediction for current MLLMs. The Sec. F in supplementary materials shows the detailed benchmark on all 13 question types.
Benchmark Examples and Predictions. Fig. 6 shows benchmark examples and qualitative results of the three most representative models that performed best on our benchmark. The top-left illustrates a sample from the Prominence field. A human can easily identify the person closest to the center as the most salient. However, both GPT-4o and Qwen2 [42] fail in this case. We include the ground truth saliency rank based on natural viewing by human observers for reference. The top-right corner shows a sample from the Subitizing field. Here, while GPT-4o and LLaVA-OV [23] fail to predict the correct quantity, Qwen2 [42]’s prediction can be automatically standardized to match the correct result. We also provide the GT salient instances, annotated and verified by humans. The bottom-left corner features a sample from the Prioritizing field, where all three models fail. The GT saliency rank represents the priority order from white to black. The bottom-right corner displays a sample from the Free-Viewing field. Option A reflects a clear pattern of human attention, while the human scanpath shown in option B is unrelated to the image. Despite this, all models incorrectly selected B. Refer to Sec. D in supplementary for additional qualitative sample questions and MLLMs predictions on all 13 question types.
Hard Case. The bottom-right shows a hard case in our HVSBench where MLLMs struggle to choose the human free-viewing scanpath, highlighting that their behavior in purely vision tasks differs from HVS. This may be because they are optimized for object recognition, not for simulating human-like viewing behaviors.
MLLMs’ Explanation on its Choice. To understand why GPT-4o errs on the prioritization task (bottom left) and to identify its focus of attention, we asked it to explain its reasoning. Specifically, GPT-4o considers: The van is large and occupies a significant portion of the image. Its size and central positioning make it visually dominant. The person near the van is secondary in prominence, as they are close to the van but are relatively smaller in size. The person standing on the right is smaller and less centrally located, reducing its visual dominance. The person on the far right is further isolated, diminishing their salience. Based on size, position, and visual weight, the van appears to be the most salient object in this image.
It highlights an over-reliance on physical attributes such as object size and position, overly associating these factors with salience. It also reveals a limitation in GPT-4o’s attention mechanism: it did not account for social or human-centric cues, which might elevate the prominence of smaller yet contextually significant objects or people.

### v2

@@ Abstract
11footnotetext: Equal contributions. Project page: https://jiaying.link/HVSBench/
Multimodal Large Language Models (MLLMs) have achieved impressive results on various vision tasks, leveraging recent advancements in large language models. However, a critical question remains unaddressed: do MLLMs perceive visual information similarly to humans? Current benchmarks lack the ability to evaluate MLLMs from this perspective. To address this challenge, we introduce HVSBench, a large-scale benchmark designed to assess the alignment between MLLMs and the human visual system (HVS) on fundamental vision tasks that mirror human vision. HVSBench curated over 85K multimodal samples, spanning 13 categories and 5 fields in HVS, including Prominence, Subitizing, Prioritizing, Free-Viewing, and Searching. Extensive experiments demonstrate the effectiveness of our benchmark in providing a comprehensive evaluation of MLLMs. Specifically, we evaluate 13 MLLMs, revealing that even the best models show significant room for improvement, with most achieving only moderate results. Our experiments reveal that HVSBench presents a new and significant challenge for cutting-edge MLLMs. Diverse human participants attained strong performance, significantly outperforming MLLMs, which further underscores the benchmark’s high quality. We believe that HVSBench will facilitate research on human-aligned and explainable MLLMs, marking a key step in understanding how MLLMs perceive and process visual information.
@@
Figure 3: We are the first to systematically study and assess MLLMs-HVS alignment, introducing diverse HVS-related questions. (a) We propose large-scale and comprehensive HVSBench, with a robust evaluation protocol. (b) We evaluate 13 leading MLLMs with HVSBench across 5 fields, revealing room for improvement and insights for developing HVS-aligned MLLMs. The shown predictions are from GPT-4o. For brevity, only 2 of 10 human scanpaths are shown. Questions and predictions are simplified. Red boxes are NOT in original image.
@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) have shown remarkable progress, achieving impressive performance across diverse vision-language tasks. These include applications such as image captioning [6], visual question answering [19, 36], document understanding [43], mathematic reasoning [35], and more. Such achievements highlight the capabilities of MLLMs in visual perception and vision-language interaction.
@@
Despite the impressive performance of MLLMs on vision tasks, we have limited understanding of why they perform well. It has been demonstrated in previous studies that principles inspired by the HVS and grounded in cognitive science [13], play a vital role in enhancing the performance of backbone models, such as attention-based architectures [51]. However, discrepancies remain in how MLLMs and humans perceive visual information. Human attention is based on innate and learned saliency, while MLLMs often perceive images as arrays of pixel values or feature embeddings. Human visual attention is sequential, adjusting based on context and prior knowledge, while MLLMs process input statically or through fixed-length attention. Human attention can be dynamically influenced and guided by goals, while MLLMs lack the cognition and the ability to “refocus” dynamically, relying purely on trained associations. Given the benefits and importance of HVS-aligned designs, including improving QA and captioning, content generation, downstream tasks, and practical applications, etc., as discussed in Sec. 4.4 , the question of whether and to what extent existing MLLMs align with HVS remains a critical area for research, especially considering the current disparities.
@@
This brings us to a fundamental question: Do MLLMs perceive the world in the same way as humans do? More specifically, do MLLMs fixate on similar regions of interest within an image or follow a similar temporal order as the HVS when perceiving an image? Humans can easily identify objects that capture their attention and perform visual searches based on context. However, in real-world scenarios, MLLMs often perceive visual information differently. For instance, in the first image of Fig. 3, if we ask, “Which object is the most visually prominent?”, most people would choose the center giraffe, while MLLMs identify the right one. Similarly, when searching for a chair in the second image, humans tend to first identify related objects (e.g., a table) and use contextual cues to help locate the chair (e.g., chairs are often near tables), while MLLMs may point to irrelevant areas. However, there has been limited research and challenges on how current MLLMs align with the HVS. On the one hand, existing public vision-language datasets are primarily designed to assess model performance on specific tasks, offering little insight into its alignment with HVS. On the other hand, traditional HVS research has largely focused on low-level, vision-only domains [38], often evaluating models based on masks or heatmaps, making it difficult to assess MLLMs where the primary output is text rather than visual data.
@@
We thoroughly evaluate 13 well-known SOTA MLLMs and human performance on HVSBench, spanning diverse architectures and model scales. This not only provides a direct comparison among these models across multiple aspects of the HVS, but also highlights the significant gap between current MLLMs and humans. Furthermore, our findings reveal critical insights for future improvements: aligning MLLMs with the HVS cannot be achieved merely by incorporating external knowledge of related cues and priors, or by integrating human-generated captions and summaries. Additionally, we stress the value of HVS alignment across domains and their evidences: mimicking human fixation enhances QA/captioning; more intuitive content generation demonstrated by our designed prominence enhancement; improved performance in HVS-specific tasks; and applications in autonomous driving and assistive tools. Our contributions are summarized as follows:
@@
    We conduct a comprehensive evaluation of 13 popular MLLMs and human using HVSBench. Additionally, we provide new insights and techniques for developing more HVS-aligned and explainable MLLMs, including a prominence enhancement method that shows how HVS-aligned MLLMs improve downstream applications.
@@ 2 Related Work
Existing benchmarks for MLLMs focus on assessing their capabilities in understanding and reasoning across modalities. A wide range of benchmarks, such as MMBench [33] and SEEDBench [27], evaluates general multimodal capabilities, including visual perception, reasoning, and comprehension. These benchmarks span diverse tasks, including document understanding [39], visual question answering [37], hallucination detection [31] and mathematical reasoning [35]. While existing benchmarks assess various abilities, they do not fully evaluate the alignment of MLLMs with the HVS. The introduction of HVSBench addresses this gap, focusing on human-centric evaluation to reveal the secret of how MLLMs perceive and process visual information. More related work on different areas (e.g., HVS and MLLM) is in the supplementary material.
@@ 3 HVSBench
In this section, we present our HVSBench in detail. We begin by defining the five distinct fields of HVS, followed by a description of multimodal QA generation, and automatic standardization and evaluation protocol. Fig. 4 shows some examples in our HVSBench. The Sec. 1 and Sec. 4 in supplementary materials provide the rationale behind selecting fields, their corresponding source datasets, additional sample questions and the results for each question type.
@@ 3.1 Preliminaries
Subitizing. Subitizing [60] is to quickly and accurately perceive the number of visually prominent objects in a scene. Unlike Prominence, it requires simultaneous attention to multiple items, vital for tasks requiring quick estimation of prominent items’ number, e.g., navigation, crowd analysis.
Prioritizing. Prioritizing in HVS ranks objects within a scene by perceptual saliency [12]. It better captures the dynamic nature of HVS, i.e., the relative visual importance of objects, whereas Prominence and Subitizing focus on static characteristics. This field has broad applications, like autonomous driving, which requires understanding relative saliency for explainable, HVS-driven decisions.
Free-Viewing. Free-viewing is crucial for both behavioral vision in the HVS and computer vision applications. It focuses on the sequence of human gaze behavior in a task-free context [55]. This behavior is bottom-up, driven solely by intrinsic visual processing. As free-viewing closely mirrors human gaze behavior, evaluating a model’s ability to predict both the order and duration of fixations, as well as a full scanpath, reveals its alignment with the HVS.
@@
The selection of the five key fields are grounded in the dual-process theory introduced in [3] and grounded by cognitive science in the introduction section of [10] of HVS, which distinguishes between bottom-up (stimulus-driven) and top-down (goal-directed) processes. Bottom-up processes mainly include prominence [20], free-viewing [21] and subitizing [50]. Top-down processes mainly include prioritizing [17] and searching [15]. These fields were chosen because they span the majority of human visual behavior, as shown in foundational reviews on visual scene processing [18], which emphasizes saliency (Prominence), attention dynamics (Free-Viewing), and search (Searching) as critical components of HVS. While others may exist, these five fields are the most widely studied and theoretically validated in HVS to the best of our knowledge.
@@ 3.2 Benchmark Curation
Variations in perception.We leverages GT human data in peer-reviewed cognitive science datasets [7, 12, 38, 54] rigorously designed for diversity and bias control. For example, Fig. 3’s source [12] employs 8 annotators (a common sample size [30] in saliency research) in diverse culture (EU/CN), race (Black/White/Asian), and gender (4M/4F), collected by strict protocols (e.g., eye-tracking monitor) to minimize bias (Sec. 3 of [12]). This ensures our benchmark captures shared human attention tendencies, aligning with common practices in HVS research.
(1) Template-Based Construction: For most questions, we generate answer options directly from GT human annotations. For example, in fields like Prominence, the options are derived from bounding boxes or other directly annotated features. To ensure diversity in the questions, we design multiple templates for each question type, following the methodology outlined in [32]. These templates provide variation while maintaining the focus of the field. For instance, in Prominence, we use templates such as:
@@ 3.3 Automatic Standardization and Evaluation
For the Prominence field, the automatic standardization cleans and normalizes predictions into consistent main choice labels (e.g., ‘A’, ‘C’, etc.) from various input formats, using regex to detect explicit or implied labels in diverse phrasing. For Subitizing, the automatic standardization processes predictions to extract integer values, including direct numbers, text-based numbers (e.g., “three”), or implied quantities. When no valid number is found, it defaults to the average ground truth value in 3.2116. For Prioritizing, except for choice answers that are processed the same as in the Prominence field, our automatic standardization detects and normalizes answers of sequence (sorting) via order-related phrasing through regex patterns. In Free-viewing and Searching fields, The automatic standardization involves extracting and normalizing scanpath data (X, Y for coordinates and T for durations) from diverse input formats, including text, JSON, and various phrasing styles. It also addresses irregularities like mismatched list lengths and incomplete scanpaths. Refer to Sec. 2 in supplementary for predictions and their corresponding standardized outputs in each field. We illustrate the superiority of our automatic standardization over LLM-based matching in Fig. 5. In our experiments, GPT-4 was prompted with the full question, the full prediction, and the text prompt: “Please extract the prediction with the correct format.” Refer to Sec. 3 in supplementary for detailed inputs, experimental settings, and pseudo-code for our automatic standardization.
@@
Evaluation Metrics. In the Prominence field in which are choice questions, the accuracy can be calculated in $\text{Acc}=\frac{\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i})}{N}$, where $y_{i}$ is the correct answer for the $i$-th question, $\hat{y}_{i}$ is the prediction for the $i$-th question., $\mathbb{I}(\hat{y}_{i}=y_{i})$ is the indicator function that equals 1 if $\hat{y}_{i}=y_{i}$ and 0 otherwise, and $N$ is total number of predictions. For the Subitizing field, we adopt widely used metrics, including Mean Absolute Error (MAE), Root Mean Square Error (RMSE), as well as Exact Match Accuracy (Acc). For Prioritizing field, the answer types include single-choice and sequence (sorting). The single choices are evaluated via accuracy. To evaluate whether the predicted order of prioritized items matches the ground truth exactly, Exact Match Accuracy (Acc) is also used. In Free-viewing and Searching fields, we employ the widely adaptive MultiMatch [21] similarity to evaluate generated scanpaths. It measures scanpath similarity across multiple dimensions: shape, direction, length, and position. In our main results, we focus on the direction (”M-Dir”) and position (”M-Pos”) dimensions for simplicity. Following standard procedures [54], predictions are cropped to a maximum length of 6, and the final scores with each prediction being the mean of the MultiMatch calculated across all 10 GT scanpaths in the image.
@@ 4.1 Evaluation Details
We use different metrics for extensive evaluations. In the Subitizing, the accuracy is computed using random sampling based on the frequency distribution of subitized object counts. For MAE and RMSE, the average number of salient objects across all samples is used as the predicted value. For MultiMatch metrics (i.e., M-Dir and M-Pos), following [54], a random guess sample is a randomly selected human scanpath of a different image. Our human evaluation involves 10 participants from diverse backgrounds considering individual differences [23]. For MultiMatch, we use a commonly used oracle method [54, 41], which compares each subject’s scanpath as a prediction to others’ as ground truth and averages the results.
We select recent open-source MLLMs for validation, including MiniGPT4-v2 [62], DeepSeek-VL [34], Idefics-series [26, 25] LLaVA-Next [32], LLaVA-OneVision (LLaVA-OV) [28], mPLUG-Owl3 [58], Qwen2-VL [52], InternVL2.0 [8] and MiniCPM-V 2.6 [57]. We utilize the $\sim$7B-scale versions of these models, a widely adopted choice for balancing efficiency and effectiveness. We also include GPT4-o [1] and Gemini-1.5 Flash [45], two representative proprietary MLLMs, as key benchmarks in our evaluation. All experiments are conducted using VLMEvalKit [14] on the same platform for consistency and fairness. We allocate 10% of HVSBench for evaluation and reserve the remaining 90% for further explorations, like instruction-tuning MLLMs.
@@ 4.2 Main Results on HVSBench
Quantitative Evaluation. Table 3 presents the evaluation results on HVSBench, highlighting that current open-source and proprietary MLLMs still underperform in aligning with HVS. Diverse human participants achieved strong performance (OA: 0.74). Given individual differences in saliency and cognitive science [23], this performance is both high and reasonable, which further validates the quality of our benchmark. Humans clearly outperform both random guessing (OA: 0.28) and MLLMs (best: 0.41) in all metrics. It is worth noting that MiniGPT4-v2, an early MLLM, has a bad performance, even worse than a random guess. This is because it often generates invalid outputs (refusing to answer or irrelevant responses) instead of outputting an option. Taking the outperforming MLLMs as a different kind of examples, GPT4-o falls short in several metrics, such as Prominence, Prioritizing, and Searching accuracy, where it performs worse than several open-source models such as Qwen2 [52]. Qwen2 [52] achieves the best overall performance, with LLaVA-OV [28] following closely behind. Both models even surpass GPT4-o overall, demonstrating the potential of recent open-source MLLMs.
@@
However, almost all methods perform poorly on scanpath similarity metrics (“M-Dir” and “M-Pos” in Free-viewing (Q9) and Searching (Q12)), underperforming random guessing, except for Qwen2 [52] in searching. These results suggest significant room for scanpath prediction for current MLLMs. The Sec. 6 in supplementary materials shows the detailed benchmark on all 13 question types.
Benchmark Examples and Predictions. Fig. 6 shows benchmark examples and qualitative results of the three most representative models that performed best on our benchmark. We include the ground truth saliency information based on natural viewing by human observers for reference. The top-left illustrates a sample from the Prominence field. A human can easily identify the person closest to the center as the most salient. However, both GPT-4o and Qwen2 [52] fail in this case. The top-right corner shows a sample from the Subitizing field. Here, while GPT-4o and LLaVA-OV [28] fail to predict the correct quantity, Qwen2 [52]’s prediction can be automatically standardized to match the correct result. The bottom-left corner features a sample from the Prioritizing field, where all three models fail. The bottom-right corner displays a sample from the Free-Viewing field. Option A reflects a clear pattern of human attention, while the human scanpath shown in option B is unrelated to the image. Despite this, all models incorrectly selected B. The bottom-right shows a hard case in our HVSBench where MLLMs struggle to choose the human free-viewing scanpath, highlighting that their behavior in purely vision tasks differs from HVS. This may be because they are optimized for object recognition, not for simulating human-like viewing behaviors. Refer to Sec. D in the supplementary for more qualitative results.
MLLMs’ Explanation on its Choice. To understand why GPT-4o errs on the prioritization task (bottom left) and to identify its focus of attention, we asked GPT-4o to explain its reasoning: The van is large and occupies a significant portion of the image. Its size and central positioning make it visually dominant. The person near the van is secondary in prominence, as they are close to the van but are relatively smaller in size. The person standing on the right is smaller and less centrally located, reducing its visual dominance. The person on the far right is further isolated, diminishing their salience. Based on size, position, and visual weight, the van appears to be the most salient object in this image. It highlights an over-reliance on physical attributes like size and position, overly associating these with salience. It also reveals a limitation in GPT-4o’s attention mechanism: neglecting social or human-centric cues, which might prioritize smaller yet contextually significant objects or people.